OpenStudy (anonymous):

3x - 2(2x - 5) = 2(x + 3) - 8 algebra. just help me pass my senior year so i can get it over with. please. i need the answer...

OpenStudy (callisto):

3x - 2(2x - 5) = 2(x + 3) - 8 Expand the terms.. 3x -2(2x) -2(-5) = 2(x) +2(3) -8 3x -4x +10 = 2x +6 -8 Can you work from here?

OpenStudy (anonymous):

i really just want the answer.... but okay. i'm really not trying to learn this. i don't need it in real life. i just need to pass algebra 2 and move on with my life...

OpenStudy (callisto):

But if you don't try , how can you get a pass?

OpenStudy (anonymous):

okay i'm not in actual school so i do this all online. it doesnt matter if i cheat. its my senior year.... i just need to get out of this and not worry about it anymore...

OpenStudy (anonymous):

ugh. guess i have to do it myself. this sucks...

OpenStudy (callisto):

Please read the Code of Conduct: '' OpenStudy values the Learning process - not the ‘Give you an answer’ process ''

OpenStudy (callisto):

You cannot figure the answer out if you don't try, perhaps this time, yes, but next time, you'll need someone to feed you with an answer. And you can never predict the future. maybe it helps you in the future

OpenStudy (anonymous):

3x-292x-5)=2(x+3)-8 3x-4x+10=2x+6-8 3x-4x+10=2x+6-8 -x+10=2x+-2 -x=2x+-12 12=3x x=12/3 x=4